# Beware of CaBER: Filament thinning rheometry doesn’t give ‘the’ relaxation time of polymer solutions

### v3

@@ 5 Conclusions and discussion
We have shown experimentally that the thinning rate of filaments of various polymer solutions is not just a material property but depends on the size of the system in CaBER with both slow (stepwise) and fast (step-strain) plate separation protocols as well as in DoS experiments, consistent with previous observations for dripping experiments [13]. Although all filaments are observed to thin exponentially, as predicted by the Oldroyd-B model (see equation 1), we show that, for CaBER with slow stepwise plate separation, the inferred apparent relaxation time $\tau_{e}$ increases with the minimum bridge radius $h_{0}$ marking the onset of capillary thinning, which is an increasing function of both the plate diameter and droplet volume, and that $\tau_{e}$ saturates at large $h_{0}$ values corresponding to plate diameters $>10$ mm significantly larger than typical CaBER plates. These observations hence suggest that CaBER relaxation times reported in the literature are not universal since testing a given polymer solution with different plate diameters and droplet volumes can yield significantly different results.
@@
We demonstrate that the variation of $\tau_{e}$ with $h_{0}$ is not caused by solvent evaporation or polymer degradation and cannot be universally explained by finite extensibility effects described by the FENE-P model. These observations suggest that the single-mode Oldroyd-B and FENE-P models miss some important features of polymer dynamics in extensional flows. The FENE-P model could only explain the variation of $\tau_{e}$ for the most dilute solution with the most viscous solvent, which is consistent with the fact that (i) the FENE-P model is derived for dilute solutions and that (ii) inertio-capillary oscillations are absent for this solution. However, since the value of the finite-extensibility parameter was chosen to optimise the agreement with experiments, we do not exclude that this agreement may also be a coincidence.
@@ Appendix B Step-strain CaBER
We used the PEOaq and PEOvisc of figure 2(b) that we label here PEOaq,old and PEOvisc,old since they were about 7 months old by the time we tested them on the Haake CaBER-1 rheometer compared to when they were tested with the slow stepwise plate separation protocol. The time evolution of the mid-filament radius $h_{mid}$ measured by the laser micrometer is shown in figures 7(a) for the PEOaq,old solution and in figure 7(b) for the PEOvisc,old solution for all plate diameters. The relaxation time $\tau_{e}$ inferred from the exponential part of the thinning dynamics, calculated from the filament thinning rate $|\dot{h}/h|$ defined as $1/3\tau_{e}$, is plotted against the plate diameter in figure 7(c) for both liquids. We find that $\tau_{e}$ increases significantly as the plate diameter increases for the PEOvisc,old solution. For the PEOaq,old solution, $\tau_{e}$ increases by a factor 2 for $2R_{0}$ between $2$ and $7$ mm and reaches a plateau for $2R_{0}\geq 7$ mm. This initial increase is not caused by experimental error, as indicated by the small error bars estimated by repeating the experiment three times for each plate. These results confirm that the apparent relaxation time increases with the plate diameter regardless of whether the plates are separated slowly or rapidly in CaBER. The range of plate diameters $3-8$ mm used with the Haake CaBER-1 rheometer is shown in grey in figure 7(c) to emphasise that, generally speaking, there is no reason to consider any apparent relaxation time measured in this range as ‘the’ relaxation time.

### v4

@@ 5 Conclusions and discussion
We have shown experimentally that the thinning rate of filaments of various polymer solutions is not necessarily just a material property but may depend on the size of the system in CaBER with both slow (stepwise) and fast (step-strain) plate separation protocols as well as in DoS experiments, consistent with previous observations for dripping experiments [13]. Although all filaments are observed to thin exponentially, as predicted by the Oldroyd-B model (see equation 1), we show that, for CaBER with slow stepwise plate separation, the inferred apparent relaxation time $\tau_{e}$ increases with the minimum bridge radius $h_{0}$ marking the onset of capillary thinning, which is an increasing function of both the plate diameter and droplet volume, and that $\tau_{e}$ saturates at large $h_{0}$ values corresponding to plate diameters $>10$ mm significantly larger than typical CaBER plates. These observations hence suggest that CaBER relaxation times reported in the literature are not universal since testing a given polymer solution with different plate diameters and droplet volumes can yield significantly different results.
@@
We demonstrate that the variation of $\tau_{e}$ with $h_{0}$ is not caused by solvent evaporation or polymer degradation and cannot be universally explained by finite extensibility effects described by the FENE-P model. These observations suggest that the single-mode Oldroyd-B and FENE-P models miss some important features of polymer dynamics in extensional flows. The FENE-P model could only explain the variation of $\tau_{e}$ for the most dilute solution with the most viscous solvent, which is consistent with the fact that (i) the FENE-P model is derived for dilute solutions and that (ii) inertio-capillary oscillations are absent for this solution. However, since the value of the finite-extensibility parameter $L^{2}$ was chosen to optimise the agreement with experiments, we do not exclude that this agreement may also be a coincidence, although this value in agrees with the microscopic prediction (equation 6).
@@ Appendix B Step-strain CaBER
We used the PEOaq and PEOvisc solutions of figure 2(b) that we label here PEOaq,old and PEOvisc,old since they were about 7 months old by the time we tested them on the Haake CaBER-1 rheometer compared to when they were tested with the slow stepwise plate separation protocol. The time evolution of the mid-filament radius $h_{mid}$ measured by the laser micrometer is shown in figures 7(a) for the PEOaq,old solution and in figure 7(b) for the PEOvisc,old solution for all plate diameters. The relaxation time $\tau_{e}$ inferred from the exponential part of the thinning dynamics, calculated from the filament thinning rate $|\dot{h}/h|$ defined as $1/3\tau_{e}$, is plotted against the plate diameter in figure 7(c) for both liquids. We find that $\tau_{e}$ increases significantly as the plate diameter increases for the PEOvisc,old solution. For the PEOaq,old solution, $\tau_{e}$ increases by a factor 2 for $2R_{0}$ between $2$ and $7$ mm and reaches a plateau for $2R_{0}\geq 7$ mm. This initial increase is not caused by experimental error, as indicated by the small error bars estimated by repeating the experiment three times for each plate. These results confirm that the apparent relaxation time increases with the plate diameter regardless of whether the plates are separated slowly or rapidly in CaBER. The range of plate diameters $3-8$ mm used with the Haake CaBER-1 rheometer is shown in grey in figure 7(c) to emphasise that, generally speaking, there is no reason to consider any apparent relaxation time measured in this range as ‘the’ relaxation time.